OpenStudy (anonymous):

(1/R)=(1/r)+(1/S) find (dR/dr) if s is a constant?

OpenStudy (anonymous):

I think i got it: 1/R = 1/r+1/s ---> 1/R = (s+r)/r*s ---> R = r*s/(s+r) So you derive it, you take out the constant s and... use the rule for deriving f(x)/g(x) d(r/r+x) = [ (r+s)(1) - r(1) ] / [r+s]^2 It all ends being( s^2) / [r+s]^2 What do you think?